OpenStudy (anonymous):

find the inverse of y=2^x-1

OpenStudy (anonymous):

y+1=x^2-1+1 =>y+1=x^2 => (y+1)^1/2=x

OpenStudy (anonymous):

ohh..i messed it up..!

OpenStudy (anonymous):

No problem, just redo, I will delete my comment for you :)